# Collectively enhanced ground-state cooling in subwavelength atomic arrays

### v2

@@ IV.2 Array cooling
Crucially, the cooling dynamics of each individual atom are determined by the properties of the collective spin states, particularly of the subradiant state, such that $\mathcal{R}_{jj}^{-}\approx\mathcal{R}_{A}^{-}/2$ [Fig. 3.(d)]. This leads to a lower average phonon number compared to the non-interacting case over a broad range of average trap frequencies $\bar{\nu}$, as shown in Fig. 4(a). To quantify the enhancement in cooling performance, we compute $\bar{n}/\bar{n}_{\text{ind}}$ and identify three distinct scenarios depending on the magnitude of $\bar{\nu}$.
@@
While the phonon number remains lower than for non-interacting atoms, it exceeds unity—implying that ground state-cooling is not achievable.
In the regime defined by Eq. (32), the steady-state phonon number scales with the subradiant decay rate $\gamma_{A}$ across all values of $\bar{\nu}$, enabling cooling of all atoms to temperatures lower than those achievable with non-interacting atoms. We thus refer to it as the array cooling regime [see central blue region of Fig. 3(e)]. As shown in Fig. 4(a,i-iii), the predictions of Eqs. (34)-(37) are in good agreement with numerical results obtained from the effective master equation (16) (solid lines), as well as from the full model including the spin dynamics (diamond markers) across a range spacings $d$ and trap frequencies $\bar{\nu}$.
@@
Stronger driving fields $\Omega$ result in faster cooling. However, for $\eta$ fixed, $\Omega$ is constrained by either Eq. (32)—which ensures the system to remain within the array cooling regime—or the weak spin-motion coupling condition $\eta\Omega\ll\gamma_{A}$. Violating either condition leads to higher steady-state phonon numbers.
To quantify these limits, we numerically solve the full coupled spin-motion dynamics Eq. (7) and extract the cooling rate $\Gamma$ by fitting the time evolution of the average phonon number $\bar{n}$. We define the critical cooling rate $\Gamma_{c}$ as the value of $\Gamma$ obtained at the maximum drive strength $\Omega$ that results in a steady-state phonon number at most 10% larger than the optimal prediction from Eq. (16) [see white line in Fig. 3(e)]. These results are shown in Fig. 4(b) for $\bar{\nu}=20\gamma_{0}$. For small trap frequency differences $\delta\nu\ll\gamma_{A}/2$, the maximum $\Omega$ is limited by the array cooling condition in Eq. (32), hence $\Gamma_{c}\propto\delta\nu$. In this case, the cooling rate is smaller than for non-interacting atoms. As $\delta\nu$ increases, the dominant constraint on $\Omega$ becomes the weak-coupling condition, and $\Gamma_{c}$ saturates to values comparable to those of non-interacting atoms. If the frequency difference becomes too large, $\delta\nu\sim\gamma_{A}/2$, the system enters the single-atom cooling regime. In this case, it is no longer possible to find a drive strength that yields an average phonon number close to the optimal prediction, resulting in the absence of markers in Fig. 4(b) at large $\delta\nu$. Figure 4(c) shows the results for $\bar{\nu}=0.25\gamma_{0}$, where motional sidebands are only resolved via the collective subradiant mode [see regime (ii) above]. In this scenario, the subradiant collective spin state enables significantly lower steady-state phonon numbers and faster cooling rates than in the non-interacting case, demonstrating a clear advantage of collective interactions. Additional detail on the numerical optimization procedure used to obtain the critical cooling rate are given in Appendix F.
@@ IV.3 Single-atom cooling
This regime is of importance if one aims at cooling a single atom using the surrounding atoms as coolant to reach lower final temperature for the target atoms. We will consider such a situation below for the case of an atom surrounded by a ring of emitters.
@@ V Collective sideband cooling of atomic arrays
Figure 5: Sideband cooling of a chain of $N$ interacting atoms. (a) Steady state phonon number per atom $\bar{n}$ and (b) critical cooling rate $\Gamma_{c}$ as a function of $N$ for various spacings $d$. The trap frequencies are normally distributed with $\bar{\nu}=20\gamma_{0}$ and $\sigma=10^{-3}\gamma_{0}$. The drive is resonant with the subradiant spin mode resulting in the smallest $\bar{n}$. The cooling rates for $d=0.2\lambda_{0}$ and $d=0.3\lambda_{0}$ exhibit similar trends and values. (c-d) Analogous for $\bar{\nu}=0.25\gamma_{0}$. In all panels, we plot the median (markers), as well as the $25$-th to $75$-th percentile region (shaded area) obtained by averaging over $300$ realizations. Diamond markers and empty circles correspond respectively to the solution of Eq. (7) and of Eq. (16). The grey dashed lines correspond to the case of non-interacting atoms $d\gg\lambda_{0}$. The solid lines are included as guides to the eye. For small systems of up to seven atoms, the steady-state phonon numbers obtained by solving the full system agree well with those predicted by the effective model in Eq. (16) (data not shown).
@@ V.1 Cooling an $N$-atom array
$|\nu_{i}-\nu_{j}|\gg\frac{2\eta^{2}\Omega^{2}}{\gamma_{\lambda}}\quad\,\forall i,j,\lambda.$ (41)
@@
To illustrate this point, we assume the trap frequencies $\nu_{j}$ are normally distributed with mean $\bar{\nu}$ and variance $\sigma^{2}$, and we address the cooling sideband that yields the lowest steady-state phonon number  [^4] For a subradiant spin state with collective linewidth $\gamma_{\lambda}$ and energy shift $J_{\lambda}$, the optimal detuning of the laser drive is $\Delta=-\sqrt{\bar{\nu}^{2}+(\gamma_{\lambda}/2)^{2}}+J_{\lambda}$.. Figures 5(a) and (c) show enhanced cooling for subwavelength chains with up to $N=30$ atoms, in both the resolved ($\bar{\nu}=20\gamma_{0}$) and unresolved ($\bar{\nu}=0.25\gamma_{0}$) regimes (analogous results hold for other geometries, see Appendix G). As in the two-atom case, the enhancement is more pronounced in the regime where the collective subradiant sidebands are resolved but the individual atomic sidebands are not. Notably, we identify conditions under which the phonon number is suppressed by more than two orders of magnitude. As $N$ increases, the steady-state phonon number begins to rise. This can be attributed to two main effects. First, the size of the array cooling region shrinks, since it becomes increasingly difficult to satisfy the condition in Eq. (41) for all frequency differences $|\nu_{i}-\nu_{j}|$. Second, while only a few collective spin states significantly contribute to cooling, a larger fraction contributes to heating.
In Figs. 5(b,d), we show the optimal cooling rate $\Gamma_{c}$, defined analogously to the two-atom case. We observe that $\Gamma_{c}$ decreases with increasing atom number, regardless of the value of $\bar{\nu}$. We attribute this behavior to the narrowing of the smallest decay rates for a few atoms  [^5] For larger number of atoms the subradiant linewidth do not scale with the system size as arguend in II, see Eq. (15)., as well as to the fact that only a few (delocalized subradiant) spin states can significantly contribute to cooling due to the collective energy shifts.
@@
Figure 6: Protocol for sequential cooling. (a) Distant blocks of a few atoms (orange boxes) are cooled simultaneously, while the spin transition of the remaining atoms (grey boxes) is detuned by and additional shift $\tilde{\Delta}\gg\gamma_{0}$. Blocks of resonant atoms that are sufficiently far apart ($L\gtrsim\lambda_{0}$) are cooled independently. (b) Cooling of a four-atom chain with spacing $d=0.15\lambda_{0}$, where the last two atoms are far detuned by $\tilde{\Delta}=-40\gamma_{0}$ (i. e., $L=d$). Light on resonance with the cooling sideband associated to the first two atoms (orange trace) leads to cooling with the same properties as for a chain of two atoms (black dashed line). The phonon number of the detuned atoms remains constant (grey line). (c) Steady state phonon number $\bar{n}$ versus cooling rate $\Gamma$ for four atoms located along the $x$-axis at positions $x_{1}=0$, $x_{2}=d$, $x_{3}=L+d$ and $x_{4}=L+2d$ for $d=0.15\lambda_{0}$. The different pairs of $\bar{n}$ and $\Gamma$ are obtained by scanning the drive strength $\Omega$. The cooling properties of a two-atom chain (dashed lines) is recovered at large $L$. We consider $\eta=0.02$ and $\nu_{j}=\bar{\nu}+\delta\nu(j-3)$ with $\bar{\nu}=20\gamma_{0}$ and $\delta\nu=10^{-3}\gamma_{0}$.
@@
V.2 Single atom cooling of an emitter in a ring
@@
We consider the symmetric scenario of Fig. 7(a), where the target atom is at the center of a ring of $N$ auxiliary atoms. For motion perpendicular to the ring and an atomic polarization that preserves the symmetry (e. g., circular in-plane or perpendicular), the target atom couples only to the symmetric ring spin mode |$\ket{S}=\sum_{n=1}^{N}\hat{\sigma}_{n}^{\dagger}\ket{0}/\sqrt{N}\equiv\hat{\sigma}_{S}^{\dagger}\ket{0}$, with collective frequency shift $J_{S}=\sum_{n=2}^{N}J_{1n}$ and decay rate $\Gamma_{S}=\sum_{n=1}^{N}\Gamma_{1n}$. Then, the cooling and heating transition rates $\mathcal{R}_{tt}^{\pm}$ only depend on the eigenstates of the non-Hermitian spin Hamiltonian
@@
where $\Delta=\omega_{L}-\omega_{0}$ is the laser detuning, and $\delta_{tS}$ captures any additional frequency offset between the target and auxiliary atoms. The coherent and dissipative couplings between the target emitter and the symmetric ring mode are $J_{tS}=\sqrt{N}J_{t1}$ and $\Gamma_{tS}=\sqrt{N}\Gamma_{t1}$, respectively. For subwavelength interatomic separation between the ring atoms, Eq. (V.2) exhibit a collective dark eigenstate |$\ket{\phi_{d}}$ with a narrowed linewidth $\gamma_{d}$, whose red-sidebands can be exploted for cooling the target atom.
Figure 7: Collective cooling of a single atom. (a) Ground state cooling of a moving quantum emitter (in free space or embedded in a mechanical resonator) via the light-induced dipole-dipole interactions with a nearby ring of $N$ emitters and radius $R$. (b) Steady state phonon number of the target resonator as a function of radius $R$ and detuning $\delta_{tS}$ for a $7$-atom ring. The black dash-dotted line and white dashed line indicated respectively Eq. (44) and the detuning leading to the minimum steady state population. (c) Decay rate of the subradiant spin mode as a function of radius for $\delta_{tS}=0$ and $\delta_{tS}^{\mathrm{dark}}$. (d) Steady state phonon number of the target emitter as a function of $N$ for a ring of radius $R=\lambda_{0}/5$ for $\delta_{tS}=\delta_{tS}^{\mathrm{dark}}$. In all panels, we consider $\eta_{t}=0.02$, $\nu_{t}=50\gamma_{c}$ and drive perpendicular to the plane defined by the ring with circular in-plane polarization.
Figure 7(b) shows the steady state population $n_{t}$ as a function of the ring radius $R$ and detuning $\delta_{tS}$. In general, smaller ring size $R$ leads to lower values for $n_{t}$. However, for small rings, the final temperature of the target atoms increase if the ring atoms and the target atoms are resonant ($\delta_{tS}=0$). This occurs because the dipole shift $J_{S}$ detunes the collective symmetric mode |$\ket{S}$ from the target atom thereby suppressing the interference effect responsible for the emergence of the narrowed linewidth $\gamma_{d}$ [blue curve in Fig. 7(c)]. This limitation can be overcome by the following detuning between the target and auxiliary emitters
@@
which minimizes the subradiant decay rate and is obtained as the detuning at which the coherent and dissipative parts of the spin Hamiltonian $\hat{H}_{S}$ commute. At small $R$, this detuning reduces the subradiant decay rate, as shown by the orange curve in Fig. 7(c), leading to a heavily reduced final steady state population [black dash-dotted curve in Fig. 7(b)] which coincides with the optimal one, obtained numerically (white dashed line). In this regime, where $\Gamma_{S}\propto N\gamma_{c}$ and $\Gamma_{cS}\propto\sqrt{N}\gamma_{c}$, the strongly subradiant state can be approximated as |$\ket{\phi_{d}}\approx(\sqrt{N}\hat{\sigma}_{t}^{\dagger}-\hat{\sigma}_{S}^{\dagger})\ket{0}/\sqrt{N+1}$. The dark mode thus describes two dipole of equal magnitude emitting with opposite phase and population mostly localized at the target atom, thereby resulting in near-perfect destructive interference [49]. At larger $R$, Eq. (44) is no longer optimal, because the decay rates of the collective modes are comparable, and it is thus favorable to tune both into resonance. Notably, $\gamma_{d}$ and $\bar{n}_{t}$ remain nearly independent of atom number at fixed $R$ under $\delta_{tS}^{\mathrm{dark}}$ [^6] Note that $\delta_{tS}^{\mathrm{dark}}$ approaches zero for 9-atom rings due to the geometric properties of the Green’s function [49]. Thus, the steady state phonon number when no detuning is applied (i. e., for $\delta_{tS}^{\mathrm{dark}}=0$) is minimized for 9-atom rings., as shown in Fig. 7(d), as $\gamma_{d}/\gamma_{0}\approx 1-\beta^{2}/\alpha+O(1/N)$ with $\beta=\gamma_{1t}/\gamma_{0}$ and $\alpha\geq\beta^{2}$ defined by $\Gamma_{S}\approx\alpha N\gamma_{t}$ [49]. Increasing $N$ beyond moderate values does not significantly improve the subradiant decay rate or steady-state phonon number.
@@ VI Conclusions
We demonstrated that the modified optical response of dipole-interacting atoms in free space can be harnessed to cool atoms to their motional ground state. Thanks to the narrowed linewidth of collective subradiant states, when the atomic trap frequencies are not identical (as is often the case in experiments), it is possible to cool the entire atomic ensemble to temperatures lower than those achievable for independent (i. e., non-interacting) atoms. Notably, we find instances where ground-state cooling is only possible due to light-induced dipole-dipole interactions. We emphasize that our scheme can be implemented using simple laser configurations (for example, by uniformly driving all atoms), requiring only the adjustment of the laser detuning.
These results are obtained from an effective master equation for the motional dynamics of dipole-interacting atoms and are confirmed by extensive numerical simulations of the original model in the Lamb-Dicke approximation. The derivation of an effective mechanical model (which holds for generic arrangements of the atoms) in the presence of subradiant states overcomes the limitations of previous approaches [29] by recognizing that atomic center of mass fluctuations broaden the subradiant collective linewidth by $\sim\eta^{2}\gamma_{0}$ [35, 36], and provide insight on the mechanisms that enable enhanced cooling.
@@
Finally, we showed for the first time that the modified collective response of subwavelength arrays can be used to improve on the control of the motional degrees of freedom of the atoms. Beyond the sideband cooling scenario considered here, we anticipate that these ideas will stimulate further exploration of the correlated dynamics between internal and external degrees of freedom, with the potential to harness these correlations to improve and enhance light–matter interfaces based on subwavelength arrays.

### v3

@@ IV.2 Array cooling
Crucially, the cooling dynamics of each individual atom are determined by the properties of the collective spin states, particularly of the subradiant state, such that $\mathcal{R}_{jj}^{-}\approx\mathcal{R}_{A}^{-}/2$ [Fig. 3.(d)]. This leads to a lower average phonon number compared with the noninteracting case over a broad range of average trap frequencies $\bar{\nu}$, as shown in Fig. 4(a). To quantify the enhancement in cooling performance, we compute $\bar{n}/\bar{n}_{\text{ind}}$ and identify three distinct scenarios depending on the magnitude of $\bar{\nu}$.
@@
While the phonon number remains lower than for noninteracting atoms, it exceeds unity—implying that ground state-cooling is not achievable.
In the regime defined by Eq. (32), the steady-state phonon number scales with the subradiant decay rate $\gamma_{A}$ across all values of $\bar{\nu}$, enabling cooling of all atoms to temperatures lower than those achievable with non-interacting atoms. We thus refer to it as the array cooling regime [see central blue region of Fig. 3(e)]. As shown in Fig. 4(a,i)-4(a,iii), the predictions of Eqs. (34)–(37) are in good agreement with numerical results obtained from the effective master equation (16) (solid lines), as well as from the full model including the spin dynamics (diamond markers) across a range spacings $d$ and trap frequencies $\bar{\nu}$.
@@
Stronger driving fields $\Omega$ result in faster cooling. However, for $\eta$ fixed, $\Omega$ is constrained by either Eq. (32), –which ensures that the system to remain within the array cooling regime, –or the weak spin-motion coupling condition $\eta\Omega\ll\gamma_{A}$. Violating either condition leads to higher steady-state phonon numbers.
To quantify these limits, we numerically solve the full coupled spin-motion dynamics equation (7) and extract the cooling rate $\Gamma$ by fitting the time evolution of the average phonon number $\bar{n}$. We define the critical cooling rate $\Gamma_{c}$ as the value of $\Gamma$ obtained at the maximum drive strength $\Omega$ that results in a steady-state phonon number at most 10% larger than the optimal prediction from Eq. (16) [see white line in Fig. 3(e)]. These results are shown in Fig. 4(b) for $\bar{\nu}=20\gamma_{0}$. For small trap frequency differences $\delta\nu\ll\gamma_{A}/2$, the maximum $\Omega$ is limited by the array cooling condition in Eq. (32), hence $\Gamma_{c}\propto\delta\nu$. In this case, the cooling rate is smaller than for non-interacting atoms. As $\delta\nu$ increases, the dominant constraint on $\Omega$ becomes the weak-coupling condition, and $\Gamma_{c}$ saturates to values comparable to those of non-interacting atoms. If the frequency difference becomes too large, $\delta\nu\sim\gamma_{A}/2$, the system enters the single-atom cooling regime. In this case, it is no longer possible to find a drive strength that yields an average phonon number close to the optimal prediction, resulting in the absence of markers in Fig. 4(b) at large $\delta\nu$. Figure 4(c) shows the results for $\bar{\nu}=0.25\gamma_{0}$, where motional sidebands are only resolved via the collective subradiant mode [see regime (ii) above]. In this scenario, the subradiant collective spin state enables significantly lower steady-state phonon numbers and faster cooling rates than in the noninteracting case, demonstrating a clear advantage of collective interactions. Additional detail on the numerical optimization procedure used to obtain the critical cooling rate are given in Appendix F.
@@ IV.3 Single-atom cooling
This regime is of importance if one aims at cooling a single atom using the surrounding atoms as coolant to reach lower final temperature for the target atoms. We consider such a situation below for the case of an atom surrounded by a ring of emitters.
@@ V Collective sideband cooling of atomic arrays
Figure 5: Sideband cooling of a chain of $N$ interacting atoms. (a) Steady state phonon number per atom $\bar{n}$ and (b) critical cooling rate $\Gamma_{c}$ as a function of NN for various spacings $d$. The trap frequencies are normally distributed with $\bar{\nu}=20\gamma_{0}$ and $\sigma=10^{-3}\gamma_{0}$. The drive is resonant with the subradiant spin mode resulting in the smallest $\bar{n}$. The cooling rates for $d=0.2\lambda_{0}$ and $d=0.3\lambda_{0}$ exhibit similar trends and values. (c-d) Analogous for $\bar{\nu}=0.25\gamma_{0}$. In all panels, we plot the median (markers), as well as the $25$-th to $75$-th percentile region (shaded area) obtained by averaging over $300$ realizations. Diamond markers and empty circles correspond respectively to the solution of Eq. (7) and (16). The gray dashed lines correspond to the case of non-interacting atoms $d\gg\lambda_{0}$. The solid lines are included as guides to the eye. For small systems of up to seven atoms, the steady-state phonon numbers obtained by solving the full system agree well with those predicted by the effective model in Eq. (16) (data not shown).
@@ V.1 Cooling an $N$-atom array
$|\nu_{i}-\nu_{j}|\gg\frac{2\eta_{j}^{2}\Omega^{2}}{\gamma_{\lambda}}\quad\,\forall i,j,\lambda.$ (41)
@@
To illustrate this point, we assume the trap frequencies $\nu_{j}$ are normally distributed with mean $\bar{\nu}$ and variance $\sigma^{2}$, and we address the cooling sideband that yields the lowest steady-state phonon number.  [^4] For a subradiant spin state with collective linewidth $\gamma_{\lambda}$ and energy shift $J_{\lambda}$, the optimal detuning of the laser drive is $\Delta=-\sqrt{\bar{\nu}^{2}+(\gamma_{\lambda}/2)^{2}}+J_{\lambda}$. Figures 5(a) and (c) show enhanced cooling for subwavelength chains with up to $N=30$ atoms, in both the resolved ($\bar{\nu}=20\gamma_{0}$) and unresolved ($\bar{\nu}=0.25\gamma_{0}$) regimes (analogous results hold for other geometries, see Appendix G). As in the two-atom case, the enhancement is more pronounced in the regime where the collective subradiant sidebands are resolved but the individual atomic sidebands are not. Notably, we identify conditions under which the phonon number is suppressed by more than two orders of magnitude. As $N$ increases, the steady-state phonon number begins to rise. This can be attributed to two main effects. First, the size of the array cooling region shrinks, since it becomes increasingly difficult to satisfy the condition in Eq. (41) for all frequency differences $|\nu_{i}-\nu_{j}|$. Second, while only a few collective spin states significantly contribute to cooling, a larger fraction contributes to heating.
In Figures. 5(b) and 5(d), we show the optimal cooling rate $\Gamma_{c}$, defined analogously to the two-atom case. We observe that $\Gamma_{c}$ decreases with increasing atom number, regardless of the value of $\bar{\nu}$. We attribute this behavior to the narrowing of the smallest decay rates for a few atoms,  [^5] For larger number of atoms the subradiant linewidth do not scale with the system size as arguend in II, see Eq. (15). as well as to the fact that only a few (delocalized subradiant) spin states can significantly contribute to cooling due to the collective energy shifts.
@@
Figure 6: Protocol for sequential cooling. (a) Distant blocks of a few atoms (orange boxes) are cooled simultaneously, while the spin transition of the remaining atoms (gray boxes) is detuned by and additional shift $\tilde{\Delta}\gg\gamma_{0}$. Blocks of resonant atoms that are sufficiently far apart ($L\gtrsim\lambda_{0}$) are cooled independently. (b) Cooling of a four-atom chain with spacing $d=0.15\lambda_{0}$, where the last two atoms are far detuned by $\tilde{\Delta}=-40\gamma_{0}$ (i. e., $L=d$). Light on resonance with the cooling sideband associated to the first two atoms (orange trace) leads to cooling with the same properties as for a chain of two atoms (black dashed line). The phonon number of the detuned atoms remains constant (gray line). (c) Steady state phonon number $\bar{n}$ versus cooling rate $\Gamma$ for four atoms located along the $x$-axis at positions $x_{1}=0$, $x_{2}=d$, $x_{3}=L+d$ and $x_{4}=L+2d$ for $d=0.15\lambda_{0}$. The different pairs of $\bar{n}$ and $\Gamma$ are obtained by scanning the drive strength $\Omega$. The cooling properties of a two-atom chain (dashed lines) is recovered at large $L$. We consider $\eta=0.02$ and $\nu_{j}=\bar{\nu}+\delta\nu(j-3)$ with $\bar{\nu}=20\gamma_{0}$ and $\delta\nu=10^{-3}\gamma_{0}$.
@@
V.2 Single-atom cooling of an emitter in a ring
@@
[…]
@@
[…]
Figure 7: Collective cooling of a single atom. (a) Ground state cooling of a moving quantum emitter (in free space or embedded in a mechanical resonator) via the light-induced dipole-dipole interactions with a nearby ring of $N$ emitters and radius $R$. (b) Steady state phonon number of the target resonator as a function of radius $R$ and detuning $\delta_{tS}$ for a seven-atom ring. The black dash-dotted line and white dashed line indicated respectively Eq. (44) and the detuning leading to the minimum steady-state population. (c) Decay rate of the subradiant spin mode as a function of radius for $\delta_{tS}=0$ and $\delta_{tS}^{\mathrm{dark}}$. (d) Steady state phonon number of the target emitter as a function of NN for a ring of radius $R=\lambda_{0}/5$ for $\delta_{tS}=\delta_{tS}^{\mathrm{dark}}$. In all panels, we consider $\eta_{t}=0.02$, $\nu_{t}=50\gamma_{c}$ and drive perpendicular to the plane defined by the ring with circular in-plane polarization.
[…]
@@
[…]
@@ VI Conclusions
We demonstrated that the modified optical response of dipole-interacting atoms in free space can be harnessed to cool atoms to their motional ground state. Thanks to the narrowed linewidth of collective subradiant states, when the atomic trap frequencies are not identical (as is often the case in experiments), it is possible to cool the entire atomic ensemble to temperatures lower than those achievable for independent (i. e., noninteracting) atoms. Notably, we find instances where ground-state cooling is only possible due to light-induced dipole-dipole interactions. We emphasize that our scheme can be implemented using simple laser configurations (for example, by uniformly driving all atoms), requiring only the adjustment of the laser detuning.
These results are obtained from an effective master equation for the motional dynamics of dipole-interacting atoms and are confirmed by extensive numerical simulations of the original model in the Lamb-Dicke approximation. The derivation of an effective mechanical model (which holds for generic arrangements of the atoms) in the presence of subradiant states overcomes the limitations of previous approaches [30] by recognizing that atomic center of mass fluctuations broaden the subradiant collective linewidth by $\sim\eta^{2}\gamma_{0}$ [36, 37] and provide insight on the mechanisms that enable enhanced cooling.
@@
Finally, we showed that the modified collective response of subwavelength arrays can be used to improve on the control of the motional degrees of freedom of the atoms. Beyond the sideband cooling scenario considered here, we anticipate that these ideas will stimulate further exploration of the correlated dynamics between internal and external degrees of freedom, with the potential to harness these correlations to improve and enhance light–matter interfaces based on subwavelength arrays.